### v1

# LTLCrit: A Temporal Logic-Based LLM Critic for Safe and Efficient Embodied Agents

Abstract
Large language models (LLMs) have demonstrated promise in reasoning tasks and general decision-making in static environments. In long-term planning tasks, however, errors tend to accumulate, often leading to unsafe or inefficient behavior, limiting their use in general-purpose settings. We propose a modular actor-critic architecture in which an LLM actor is guided by LTLCrit, a trajectory-level LLM critic that communicates via linear temporal logic (LTL). Our setup combines the reasoning strengths of language models with the guarantees of formal logic. The actor selects high-level actions from natural language observations, while the critic analyzes full trajectories and proposes new LTL constraints that shield the actor from future unsafe or inefficient behavior. The architecture supports both fixed, hand-specified safety constraints and adaptive, learned soft constraints that promote long-term efficiency. Our architecture is model-agnostic: any LLM-based planner can serve as the actor, and LTLCrit serves as a logic-generating wrapper. We formalize planning as graph traversal under symbolic constraints, allowing LTLCrit to analyze failed or suboptimal trajectories and generate new temporal logic rules that improve future behavior. We evaluate our system on the Minecraft diamond-mining benchmark, achieving 100% completion rates and improving efficiency compared to baseline LLM planners. Our results suggest that enabling LLMs to supervise each other through logic is a powerful and flexible paradigm for safe, generalizable decision making.
1 Introduction
Large Language Models (LLMs) have recently demonstrated strong performance on diverse reasoning and decision-making tasks, from natural language understanding, reasoning (Huang & Chang, 2022; Wei et al., 2022), and code generation (Li et al., 2022; Chen et al., 2021). However, much of this success has been in static, text-based settings. To deploy AI systems in real-world scenarios, it is essential to build agents capable of safe and efficient decision-making in complex, partially observable environments with long-term objectives and uncertainty.
Embodied environments such as Minecraft (Fan et al., 2022), Habitat (Savva et al., 2019), Alfworld (Shridhar et al., 2020) serve as useful testbeds for the next generation of agents. These environments require agents to perceive complex states, plan for long-horizon problems, under sparse supervision and rewards. Recent work has begun to explore the use of LLMs in embodied settings. SayCan Ahn et al. (2022) uses LLMs to decide affordance filtered actions for short-horizon robotic tasks, while Voyager (Wang et al., 2023a) leverages an LLM to autonomously explore and obtain as many new materials as possible in Minecraft.
Despite these advances, LLMs continue to struggle with long-term sequential planning in embodied dynamical environments. While LLMs can often generate coherent short-term plans or respond to individual prompts, they lack the consistency, memory, and iterative refinement required for solving multi-step tasks where intermediate actions must align with long-term objectives. These challenges are especially pronounced in open-ended domains like robotics or Minecraft, where the agent must choose from a large set of possible actions while reasoning about tools, environment dynamics, and partial observability.
Recent evaluations (Kambhampati et al., 2024) show that LLMs struggle to reliably solve classical planning problems without the aid of external verification. Even simple domains expose brittleness: small prompt variations can cause cascading errors, and plans often violate task preconditions or logical constraints. These limitations have spurred the development of hybrid frameworks that pair LLMs with verifiers (Silver et al., 2022; Guo et al., 2024). While these methods improve correctness, they rely on hand-written task models.
Ultimately, a central challenge is enabling foundational models like LLMs to act as reliable, autonomous agents in real-time, safety-critical, multi-agent environments such as healthcare (Hosny et al., 2018), automated transportation (Wang et al., 2023b), and domestic assistance (Birkmose et al., 2025). In each of these contexts, both safety and efficiency are crucial. Unsafe or inconsistent behavior can lead to failures in physical systems, and inefficient or incompetent agents rapidly lose the trust of human collaborators (Esterwood & Robert Jr, 2023). We argue that planning architectures should combine the reasoning strengths of LLMs with the formal structure of symbolic logic to ensure both flexibility and safety in real-world decision making.
1.1 Related works
Temporal Logic for Planning and Reinforcement learning
Linear Temporal Logic (LTL) (Pnueli, 1977) is a formal language for specifying temporal properties of systems, for example, conditions that must hold at the next timestep given the current state, or conditions that must always hold, or eventually become true. LTL has been widely used in robot motion planning (Fainekos et al., 2005), for safe planning and control (Wongpiromsarn et al., 2012), and even in reinforcement learning Alshiekh et al. (2018). In each of these applications, LTL is used to specify safety constraints; however, one may also use it to specify efficiency constraints, as will be discussed in our graph formulation.
Language Models for Planning and policy learning
Recent work has shown some promise in terms of LLMs’ abilities to perform in short-horizon planning tasks (Huang et al., 2022a). Particularly, Ahn et al. (2022) utilizes an LLM to asses the value of a given action, and weights it with an affordance function which quantifies the feasibility or success rate of the action given the current state of the system. It achieves a strong success rate in short-horizon mobile manipulation tasks. Huang et al. (2022b) introduces a closed-loop inner monologue by feeding language-based feedback back into the LLM during execution, enabling re-planning in dynamic environments. Simultaneously, other works such as Kambhampati et al. (2024) suggest that LLMs do not show reliability while operating autonomously, and require external critics. Our work builds directly upon these ideas, augmenting LLMs with formal verifiers. Finally, a majority of the work in LLM planning and policy learning talks about the success rate. Our emphasis is not only on completion rates, but also on speed of completion.
Integrating Symbolic Reasoning with LLMs
There is also a growing body of work at the intersection of LLMs and temporal logic; however, the key emphasis is on using LLMs as tools for translation between natural language and LTL constraints Chen et al. (2023); Liu et al. (2023). Another line of work uses these translators in the loop (Yang et al., 2024), enforcing these constraints to filter out unsafe actions during execution. Compared to prior literature, which enables the use of static handcrafted rules, we allow an LLM critic to come up with further LTL constraints that show empirical efficiency gains in LLM planning problems.
LLM-guided Actor-Critic Architectures
Recent work explores hybrid actor-critic setups where LLMs guide planning and evaluation, often using natural language feedback or prompt optimization loops (Dong et al., 2023; Yang et al., 2025). These models rely on soft, unstructured feedback and lack formal guarantees, particularly in multi-step embodied reasoning tasks. Our LTL-based trajectory critic offers interpretable, symbolic evaluation aligned with both safety and efficiency constraints.
1.2 Contributions
We propose LTLCrit, a novel Linear Temporal Logic-based Critic, which augments and boosts the performance of existing off-the-shelf LLM planners by protecting the planners against unsafe and inefficient actions. The composition of LTLCrit with an existing LLM planner leads to a novel symbolic actor-critic architecture designed to solve long-horizon planning tasks in dynamic, embodied environments using large language models (LLMs). Planning is decomposed across two timescales, a online actor loop that proposes high-level actions based on current state descriptions, and a offline critic loop that imposes symbolic constraints learned from past trajectories. Our modular decomposition aligns with LLMs’ strengths in local reasoning while addressing their weaknesses in long-term consistency over sequential domains.
LTLCrit is domain-agnostic, extensible to a variety of embodied tasks, and naturally integrates with existing LLM-based planners. It is the first to use symbolic temporal logic as a communication protocol between actor and critic modules, enabling interpretable, verifiable, and generalizable decision making. Our contributions are threefold.
1.
Symbolic Actor-Critic Architecture: We propose a two-timescale architecture where an LLM actor generates high-level actions in real time, and an LLM critic periodically analyzes trajectories to induce Linear Temporal Logic (LTL) constraints over symbolic abstractions of state and action. These constraints are used to prune unsafe or inefficient decisions, leading to improved sample efficiency and task performance. Our formulation allows constraints to be automatically discovered and continuously refined over time.
2.
Communication via Temporal Logic: We introduce a novel mechanism for actor-critic interaction based on symbolic temporal logic. In contrast to traditional reinforcement learning critics or natural language feedback, our critic outputs verifiable, machine-checkable LTL constraints. LTL constraints provide strong safety guarantees, enforces logical consistency, and enables constraint reuse across similar states and tasks. We also present a graph-theoretic abstraction of planning under temporal logic, which highlights the role of constraint pruning in reducing planning complexity.
3.
Empirical Validation in Embodied Settings: We evaluate our method on challenging long-horizon tasks in Minecraft, where agents must autonomously mine a diamond. Our critic can improve both the consistency of behavior and average performance for two popular LLM planners, SayCan (Ahn et al., 2022) and InnerMonologue (Huang et al., 2022b).
Together, these contributions advance the goal of safe, efficient, and interpretable LLM-based decision-making in complex, dynamic environments. Our symbolic actor-critic framework bridges the gap between natural language reasoning and formal planning, opening up new possibilities for trustworthy embodied agents.
2 Background and Problem Formulation
Enabling LLM-based agents to work effectively alongside humans in embodied environments is a fundamental goal of LLM research. This has motivated recent efforts to test and enhance LLMs’ abilities in interactive, multi-step environments with long-term objectives. A central challenge LLM agents face in such settings is the lack of goal formalization and the difficulty of performing coherent long-horizon planning. Without structured guidance, LLM agents can behave reactively, inconsistently, or in an unsafe way. These behaviors are incompatible with reliable human-robot teaming.
Moreover, beyond safety, efficiency plays a critical role in collaboration. In human-agent teams, inefficient partners are perceived as unreliable, leading humans to reduce trust, ignore suggestions, or take full control (Fan et al., 2008). To deploy LLM agents in high-stakes or collaborative domains, we must ensure that their behavior is not only safe but also efficient, interpretable, and improvable over time.
In our work, we address these challenges by developing a learning mechanism that imposes formal guarantees on the behavior of LLM agents. Specifically, we aim to design agents that can self-correct and improve efficiency through structured feedback, from both humans and LLMs, using relatively little data. We study the planning problem in the context of Minecraft, an embodied environment, making LLM agents achieve the goal of mining a diamond.
2.1 Problem Statement
We formalize the problem of safe and efficient LLM planning in embodied environments as follows. Consider an agent operating in an environment with state space and a finite high-level action space , where . The agent observes a representation of the state, which is modeled by . Unlike standard RL settings, we assume no explicit reward function, leveraging instead the zero-shot learning capabilities of LLMs via a high-level goal description (e.g., “make an iron pickaxe” or “obtain a diamond”). We enforce safety and enable systematic constraint enforcement using an additional abstract boolean representation, , where , consisting of atomic propositions capturing important environmental features.
Within , there are two sets of interest.
1.
is the set of states where the high-level objective is satisfied.
2.
is the set of states that violate safety constraints.
We enforce behavioral constraints using Linear Temporal Logic (LTL) formulas over . We have two kinds of constraints.
Safety Constraints are expert-verified, hand-authored LTL constraints that prevent transitions into . These capture fundamental requirements such as collision avoidance.
Efficiency constraints are automatically generated by LTLCrit and eliminate suboptimal behaviors observed during execution, emerging from patterns such as redundant action sequences. For the sake of simplicity, we assume that every high level action is equally expensive to execute, and our goal is to minimize the number of primitives required to achieve the goal.
We seek to develop an agent that satisfies all safety constraints with high probability and efficiently performs a sequence of actions that leads to goal states.
2.2 Linear Temporal Logic Primer
Linear Temporal Logic(LTL) (Pnueli, 1977) provides a formal language for expressing temporal properties over sequences of states. LTL formulas are built using the usual Boolean operators, and four temporal statements:
1.
: holds at the next timestep.
2.
: eventually holds.
3.
: holds in all future states.
4.
means that holds until becomes true
LTL formulas can be converted into Büchi automata, which are finite-state machines that accept infinite sequences. Büchi automata enable the algorithmic verification of whether a trajectory satisfies a given temporal specification.”. Conversion into a Büchi automata allows us to systematically validate agent actions against formal constraints before execution. LTL formulas defined over atomic propositions that have direct physical meaning are guaranteed to be human-interpretable. In our work, we utilize SPOT (Duret-Lutz & Poitrenaud, 2004) to enforce LTL laws upon the LLM actor.
2.3 Minecraft
Minecraft has emerged as a popular domain for studying embodied agents for long-horizon reasoning and open-ended planning. It is a partially observable dynamical environment, with a very large state space and action space. Its compositional task structure makes it well-suited to be a benchmark for complex planning problems. In our experiments, we aim to solve the problem of obtaining a diamond in Minecraft, inspired by (Guss et al., 2019). Mining a diamond requires the completion of a series of subgoals, with no immediate reward signal. More recent studies incorporate language and vision, such as MineDojo Fan et al. (2022), which utilizes video and text data on the internet with a programmable agent API. Another line of work (Wang et al., 2023a) discusses the use of LLMs to design lifelong learning agents based on environment feedback and skill reuse. However, none of these methods presents guarantees on safety. We interface with Minecraft via the Mineflayer API (PrismarineJS Team, 2025), which provides structured local observations and built-in path-planning utilities. The agent interacts with the world using a fixed set of primitive commands. Experiments are conducted in different worlds with random seeds to account for robustness to the map.
3 Methods
3.1 Architecture Overview
We propose a hierarchical planning framework that integrates large language models (LLMs) with formal symbolic reasoning to achieve safe and sample-efficient decision-making in complex environments. The architecture consists of two interacting loops operating at different timescales:
Online actor loop: At each timestep, an LLM actor is given a natural language state description, and chooses a high-level action via an LLM-based actor. The chosen action is verified against all current LTL constraints, which are expressed over the symbolic abstraction . If valid, the action is executed via a low-level controller. If not, the actor is informed that the chosen action is invalid and prompted to choose a new action. A few examples illustrating the operation of the LTL verifier are in Figure 2.
Offline critic loop: Periodically, an LLM-based critic module analyzes completed trajectories to identify inefficient behavior and propose new LTL constraints, and eliminate existing constraints. These changes are immediately pushed to the verifier, instantly impacting future behavior.
Our design, which decouples online, reactive decision making and offline refinement of decisions, enables safe, sample-efficient, and interpretable planning. Our modularity also allows for the ease of transfer across domains. A pictorial representation of our architecture is presented in Figure 1.
3.2 Planning as a Graph Traversal Problem
We frame efficient planning as a shortest path problem under safety constraints. We assume that each primitive action incurs a unit cost, as LLM calls are the bottleneck in terms of execution speed. The agent’s object is to reach a state in from the current state in as few steps as possible, while avoiding . One may view the planning problem in two graph-theoretic ways.
1. Semi-MDP View:
We may treat as the state space of a Markov Decision Process(MDP) with unknown dynamics. For an agent to trade off between exploring to learn transition probabilities and an efficient policy.
2. Bipartite graph View:
Modeling LTL laws in the Semi-MDP framework involves constructing a lifted product graph with additional nodes. We define a bipartite graph , whose nodes comprise two disjoint sets of symbolic states and state-action pairs . Edges from to are allowed only if is permitted by current LTL constraints in state . Edges from to represent the state transition function given a fixed action. The bipartite structure makes explicit the roles of the actor (choosing edges from to ) and the critic (pruning these edges via constraints).
In both of these views, we attempt to solve a path planning problem on a graph with nodes, and an unknown structure. As a consequence, solving the shortest path planning problem optimally is also at least . We resort to using LLMs for natural language reasoning-driven guided exploration to overcome the exponential complexity.
3.3 LLM Actor: Guided Exploration
The LLM actor receives a full state description and proposes a high-level action from the action space such as “mine_stone” or “craft_wooden_pickaxe”. In the bipartite graph , choosing a high level action given a state corresponds to selecting an outgoing edge from the current state node. Validity is enforced via a Büchi automaton, which is defined based on the LTL constraints. The use of the LLM allows us to replace brute force exploration with semantically guided traversal in a large symbolic space. Our modular nature allows us to directly utilize existing LLM planning algorithms from the literature for our actor. We adopt Saycan (Ahn et al., 2022) and InnerMonologue (Huang et al., 2022b) as actors, and also conduct some experiments on a baseline naive LLM actor.
3.4 LTLCrit: the Linear Temporal Logic-based LLM Critic
The critic operates at a slower timescale, analyzing complete trajectories to identify inefficiencies, and proposing new LTL constraints of the form
where is a boolean expression over symbolic state features, and is a boolean formula over one-hot encoded actions such that at least one valid one-hot vector in satisfies . These constraints eliminate inefficient behaviors. For example, the constraint
prevents agents from crafting duplicate wooden pickaxes. While the agent may have only visited some of the states where the agent has a wooden pickaxe, it’s reasoning capabilities allow it to make general claims about all states where the agent does has a wooden pickaxe. To regulate the critic’s behavior and ensure data-grounded generalization, we restrict it to three sources of constraint generation.
1.
Environment Feedback: When the actor attempts an invalid action (e.g., mining diamonds without an iron pickaxe), the environment throws an error. While the actor sees this error and modifies its action immediately, the critic encodes this failure into a permanent constraint, shielding future policies from repeating the mistake.
2.
Graph-Based Efficiency Analysis: Through prompting, we ask the LLM to think of the symbolic task graph as described in section 3.2. We ask the critic to classify actions as efficient and inefficient, promoting movement towards the goal states and limiting wasteful actions.
3.
Overconstrained States: When current laws collectively eliminate all feasible actions in a state (typically due to overgeneralization on the part of the critic), the system detects an overconstraint failure. In such cases, the actor falls back to a minimal set of hand-engineered safety constraints. Offline, these conflict states and the most constraining laws are recorded and analyzed by the critic, proposing refined or relaxed constraints that avoid blocking all behavior.
Crucially, the critic is only allowed to induce constraints grounded in trajectory data. That is, each must be true for at least one observed state in the trajectory, ensuring that all proposed rules are traceable.
Theorem 1.
Let be a trajectory with , and each a one-hot vector representing an action from a finite action space of size . Consider an algorithm that generates LTL constraints of the form in equation 1 where is a boolean condition over symbolic state variables that holds for at least one in the trajectory and .
Then, the algorithm can generate at most distinct such laws from the trajectory while ensuring that, for every state , there exists at least one action satisfying all LTL laws.
Proof.
The trajectory contains state-action pairs , so there are at most unique values for . Since each constraint is of the form , and must hold for at least one , the number of distinct that can be constructed from the trajectory is at most . Further, the algorithm can only decide if each action is allowed or disallowed. As a consequence, the total number of laws is . However, since actions cannot be simultaneously allowed and disallowed, the number of actions an algorithm operating under our assumptions can make such that every state has at least one feasible action is . ∎
Our bipartite graph representation of the planning problem produces a natural upper bound of possible edges from the state partition to the state action partition. Pruning the bipartite graph using only observed trajectory data drastically reduces the complexity to linear in dataset size. Furthermore, our use of interpretable natural language labeled atomic propositions allows the LLM critic to generalize constraints beyond observed states. For example, when a constraint like equation 2 is learned in one state where the LLM had a pickaxe and tried to craft another one, the LLM can reliably apply that rule to semantically equivalent but unseen states.
Finally, the sparse structure of most goal-directed tasks ensures that very few actions are efficient at each decision point. The sparsity of the underlying optimal graph, combined with the critic’s structured constraint generation and the LLM’s generalization ability, enables sample-efficient and robust agent behavior.
4 Experiments
4.1 Experimental Setup
Next, we evaluate our architecture on a long-horizon diamond mining task in Minecraft. We utilize OpenAI’s GPT 4.1 for the LLM actor, and o3-mini for the LLM critic. Temperatures are set to 0.1 to reduce stochasticity in code generation. Our environment is implemented via Mineflayer’s JavaScript API, which exposes structured access to agent state and motor control primitives. On top of this, we define a set of atomic action wrappers (for example, mineBlock, craftItem) and expose structured observations to the LLMs. Further details are discussed in Appendix A.1.
4.2 Baselines
Our modular architecture allows us to utilize off-the-shelf LLM actors and improve them using our symbolic critic. In particular, we focus on augmenting two existing LLM-based planners that represent distinct paradigms in LLM control.
SayCan(Ahn et al., 2022)
The SayCan actor consists of two stages: affordance filtering and value-guided selection. First, an affordance function decides if a high-level action is feasible given the current agent and environment state. Then, a language model assigns each filtered feasible high-level action a value estimate reflecting how helpful it is for achieving the final goal (in our case, mining a diamond). The agent then executes the feasible action with the highest predicted value. In order to implement an affordance filter in our framework, we construct the filter using a series of LTL laws. The list of these laws that form the affordance filter is in Appendix A.3.
InnerMonologue(Huang et al., 2022b)
Inner Monologue augments an LLM planner with a persistent stepwise reasoning across timesteps. At each step, the agent outputs a natural language “thought” along with a code snippet to execute, and then receives feedback from the environment. The loop is repeated, with the prompt at each round including the previous thoughts, code, and outcome. This enables implicit reflection and planning over time.
We augment both of these planners with our critic, allowing us to measure how symbolic constraints improve performance and safety. To derive the LTL constraints, we analyze the actor-only trajectories using the LLM critic, which generates candidate LTL rules that promote safe and efficient behavior. The resulting rules are added to a growing constraint set. If suboptimal behavior is observed under the current constraint set, we re-run the resulting trajectories through the critic, which may propose additional constraints or refinements. This iterative process continues until performance stabilizes. In practice, InnerMonologue required 2 iterations to stabilize. We present all relevant prompts and discovered LTL laws in the appendix.
We do not compare to vision-based Minecraft agents trained on pixel observations with continuous motor outputs (Nottingham et al., 2023; Cai et al., 2023; Wang et al., 2023c). Our goal is to improve LLM planning in environments with structured state and action spaces. Our framework assumes access to state descriptors and motor-level wrappers, which are common in many robotics pipelines. As such, our results are aligned with improving task efficiency and safety in high-level LLM-driven control.
4.3 Performance Metrics
We test LLM-based planners in a standard Minecraft environment where the agent must craft a diamond tool from scratch, involving long-horizon dependencies across mining and crafting subgoals. We evaluate our agents along two axes:
Efficiency
We record the number of primitive actions taken to reach key sub-goals, collecting wooden tools, stone tools, iron tools, and finally a diamond. As shown in Table 1, the critic consistently improves efficiency in each subgoal, particularly, the jump from stone tools to iron tools involves complex dependencies and parallel subgoals that InnerMonologue finds difficult. We note a 23% increment in the performance of InnerMonologue in the diamond mining task. SayCan without LTL is very easily distracted by the abstract nature of high level actions such as “explore”, and does not make meaningful progress on the task. LTL crit identifies these drawbacks and blocks exploration related actions till the right tools are available.
Method | Wooden Tool | Stone tool | Iron Tool | Diamond |
SayCan | N/A (0/5) | N/A (0/5) | N/A (0/5) | N/A (0/5) |
SayCan + LTLCrit | 12.6(5/5) | 17.6(5/5) | 37.8(5/5) | 45.4(5/5) |
InnerMonologue | 12.2 (5/5) | 18.2 (5/5) | 43.25 (4/5) | 45.5 (4/5) |
InnerMonologue + LTLCrit | 9.4 (5/5) | 14.4 (5/5) | 32.0 (5/5) | 35.8 (5/5) |
Safety
We log invalid or failed actions and measure the total number of failed actions and the number of unsafe actions blocked by the LTL critic. Our critic significantly reduces failure rates. Common failure modes include attempting to mine iron or diamond blocks without appropriate tools. These are consistently filtered by our LTL constraints. We note that the number of failed actions with LTLCrit is nonzero due to a combination of two factors. First, our abstracted observation variables are not expressive enough, and second, Mineflayer’s internal path planner is not perfect and returns errors sometimes. All the laws designed by LTLCrit are presented in Appendix A.3.
Method | Failed Actions | Critic-Blocked Unsafe Actions |
InnerMonologue | 23% | N/A |
InnerMonologue + LTLCrit | 4.5% | 15% |
5 Discussion
5.1 Atomic propositions are a key design choice
Atomic propositions are the variables used to build LTL statements. In our case, they are the boolean variables that are associated with and . The choice of atomic propositions is a fundamental design choice for our architecture, as it directly controls the expressivity and the precision of the LTL constraints that are accessible to the LLM. For example, our observation has a general atomic proposition for having fuel (which is used to convert raw iron into iron ingots). However, our proposition does not specify what kind of fuel, and the LLM agents often choose fuel sources they do not possess, leading to errors. Choosing interpretable atomic propositions with names based on natural language leads to better reasoning on the part of the LLM critic.
5.2 Nature of the induced constraints
Our structured process of guiding the LLM critic through LTL constraint generation encourages the formation of correct LTL laws. The LLM critic reliably identifies efficient actions near the start state and the goal state, and then works its way inwards over iterations. Our use of LTL constraints allows the rules to be interpretable, human editable, and, most importantly, verifiable. Our critic is adaptive and makes rules depending on the errors of the particular actor it is working with. For instance, the SayCan actor is easily distracted from long-term goals if the action label is similar to the long-term goal. For instance, actions describing exploration are strongly preferred by SayCan. Our critic notices this bias and constrains the actor to not exploring for diamonds until the correct prerequisites are achieved. On the other hand, InnerMonologue requires guidance at the midpoint of the task, where it often repeats basic actions over making progress towards the goal.
5.3 Prompt engineering has limited guarantees
Prompt engineering has been shown to improve zero-shot task performance. However, it does not provide guarantees of correctness, safety, or logical consistency. In safety-critical settings, relying solely on natural language prompting is insufficient. On the other hand, having a formal verifier in the loop, filtering unsafe or inefficient actions regardless of prompt quality, leads to guarantees on behavior. While our LLM actor may be improved by prompt engineering, the verifier and the critic are crucial components in any safety-critical environment.
5.4 Limitations
Despite its strengths, our approach has several important limitations. First, the quality of induced LTL constraints is tightly coupled to the symbolic abstraction: if key atomic propositions are missing or poorly defined, the critic may fail to express necessary constraints. We mitigate this with natural language–grounded APs. However, automatic discovery of meaningful abstractions is an open problem. Second, the critic operates offline and requires multiple trajectories to converge on effective rules. This limits applicability in fully online or real-time settings. Operating the critic online tends to lead to over-constraining laws. Third, although the verifier guarantees logical correctness, it does not optimize for reward; excessively conservative constraints may limit exploration or induce deadlocks. Fourth, although human editability is a strength, it introduces a new form of manual overhead: deciding when and how to intervene in the constraint set. Finally, our method assumes access to symbolic task structure, which may not always be available or easily extracted in complex environments.
6 Conclusion and Future work
We introduced a modular actor-critic architecture in which an LLM critic supervises an LLM actor using linear temporal logic (LTL) constraints over abstracted versions of full trajectories. The critic operates at a slower timescale and applies zero-shot reasoning to identify and correct unsafe or inefficient behavior in a few iterations. Our modular architecture enables us to use existing off-the-shelf LLM planners as actors. The use of LTL constraints guarantees shielding of the LLM from unsafe or inefficient behavior.
By expressing constraints in LTL over human-readable atomic propositions, we gain a symbolic structure that is immediately enforceable and human verifiable. Applied to the Minecraft diamond-mining task, our system efficiency gains over baseline LLM planners and eliminates entire classes of planning failures. Future work includes learning atomic propositions automatically, extending the critic to multi-agent coordination domains, and extending these experiments to real-life human-robot teaming contexts. Our results suggest that formal logic-based supervision is a promising path toward safe, scalable, and general-purpose LLM agents.
Acknowledgments
This research was funded by the Office of Naval Research, MURI N00014-22-1-2813. We thank Arvind R. Venkatakrishnan for his assistance in the design of figures. We acknowledge the use of ChatGPT for assistance in improving the wording and grammar of this document.
Appendix A Appendix
A.1 Interfacing with Minecraft
We interface with Minecraft using the Mineflayer API, which exposes high-level observations as structured JSON objects and provides access to built-in path-planning routines. We define a set of action primitives on top of this interface to abstract the agent’s decision space while preserving task complexity. Each primitive internally invokes Mineflayer’s planners and lower-level control routines. The full list of primitives is as follows:
1.
mineBlock(bot, blockName): Mines 1 block of type blockName, provided it is visible within a 32-block radius.
2.
placeItem(bot, blockName, position): Places a block at a specified position, assuming the location is unoccupied and adjacent to an occupied block.
3.
craftItem(bot, itemName): Crafts 1 item of type itemName, assuming all ingredients are present in the agent’s inventory. Recipes that require a crafting table assume one is nearby.
4.
smeltItem(bot, itemName, fuelName): Smelts 1 item of type itemName using fuelName, assuming both are present in the agent’s inventory and a furnace is nearby.
5.
equipItem(bot, itemName, destination): Equips tools or armor. Some blocks require a minimum tool tier to mine, and the appropriate tool must be equipped.
6.
exploreUntil(bot, direction, condition): Causes the agent to explore in a specified direction until a user-defined condition is met (e.g., locating a specific block).
These primitives allow for rich compositional behavior while delegating locomotion to Mineflayer’s planners. The full sequence of subgoals required to successfully mine a diamond in this setup is as follows:
1.
Obtain wooden logs.
2.
Craft logs into sticks and planks.
3.
Craft a wooden pickaxe.
4.
Mine stone blocks using the wooden pickaxe.
5.
Craft a stone pickaxe and a furnace.
6.
Obtain raw iron by mining iron blocks.
7.
Smelt raw iron into iron ingots.
8.
Craft an iron pickaxe.
9.
Explore the world and mine a diamond block.
This structured task allows us to evaluate the ability of our actor-critic framework to perform multi-stage reasoning, tool use, and resource management over long horizons.
A.2 List of atomic propositions
We have two sets of atomic propositions, one for observations and one for high-level actions.
Proposition | Meaning |
obs_has_log | Agent has at least 1 log |
obs_has_plank | Agent has at least 1 plank |
obs_has_2x_plank | Agent has planks |
obs_has_3x_plank | Agent has planks |
obs_has_4x_plank | Agent has planks |
obs_has_11x_plank | Agent has planks |
obs_has_2x_stick | Agent has sticks |
obs_has_3x_cobble | Agent has cobblestone |
obs_has_8x_cobble | Agent has cobblestone |
obs_has_11x_cobble | Agent has cobblestone |
obs_has_wood_pickaxe | Agent has wooden pickaxe |
obs_has_stone_pickaxe | Agent has stone pickaxe |
obs_has_iron_pickaxe | Agent has iron pickaxe |
obs_has_diamond | Agent has at least 1 diamond |
obs_has_iron_ingot | Agent has iron ingot |
obs_has_3x_iron_ingot | Agent has iron ingots |
obs_has_1x_iron_ore | Agent has iron ore |
obs_has_2x_iron_ore | Agent has iron ore |
obs_has_3x_iron_ore | Agent has iron ore |
obs_has_crafting_table | Agent has crafting table |
obs_has_furnace | Agent has furnace |
obs_has_fuel | Agent has fuel (e.g., coal) |
obs_near_crafting_table | Agent is near crafting table |
obs_near_furnace | Agent is near furnace |
obs_diamond_in_chunk | Diamonds detected nearby |
obs_iron_in_chunk | Iron ore detected nearby |
obs_coal_in_chunk | Coal detected nearby |
obs_iron_pickaxe_equipped | Iron pickaxe is equipped |
obs_stone_pickaxe_equipped | Stone pickaxe is equipped |
obs_wood_pickaxe_equipped | Wooden pickaxe is equipped |
Action | Meaning |
action_mine_log | Mine wood logs |
action_mine_stone | Mine stone |
action_mine_iron_ore | Mine iron ore |
action_mine_coal | Mine coal |
action_mine_diamond | Mine diamond |
action_craft_planks | Craft planks from logs |
action_craft_stick | Craft sticks from planks |
action_craft_wooden_pickaxe | Craft wooden pickaxe |
action_craft_stone_pickaxe | Craft stone pickaxe |
action_craft_iron_pickaxe | Craft iron pickaxe |
action_craft_crafting_table | Craft a crafting table |
action_craft_furnace | Craft a furnace |
action_smelt_iron | Smelt iron ore into ingots |
action_equip_wood_pickaxe | Equip wooden pickaxe |
action_equip_stone_pickaxe | Equip stone pickaxe |
action_equip_iron_pickaxe | Equip iron pickaxe |
action_explore_general | Explore randomly |
action_explore_diamond_down | Explore downward for diamonds |
action_place_crafting_table | Place crafting table |
action_place_furnace | Place furnace |
A.3 List of LTL laws imposed for each actor
A.3.1 SayCan
The hard-hand engineered safety rules that prevent illegal actions are given below:
1.
LTL:
Explanation:Diamonds cannot be mined unless an iron pickaxe is equipped. -
2.
LTL:
Explanation:Cannot craft any type of pickaxe unless near a crafting table. -
3.
LTL:
Explanation:Cannot craft a furnace unless near a crafting table. -
4.
LTL:
Explanation:Cannot mine iron ore unless a stone or iron pickaxe is equipped. -
5.
LTL:
Explanation:Cannot mine stone unless any pickaxe is equipped. -
6.
LTL:
Explanation:Cannot equip an iron pickaxe unless the agent has one. -
7.
LTL:
Explanation:Cannot craft a wooden pickaxe without enough planks and sticks. -
8.
LTL:
Explanation:Cannot craft a crafting table without 4 planks. -
9.
LTL:
Explanation:Cannot craft a furnace without 8 cobblestone. -
10.
LTL:
Explanation:Cannot craft an iron pickaxe without 2 sticks and 3 iron ingots. -
11.
LTL:
Explanation:Cannot craft a stone pickaxe without cobblestone and sticks. -
12.
LTL:
Explanation:Cannot mine coal unless coal is nearby and a pickaxe is equipped. -
13.
LTL:
Explanation:Cannot craft planks without logs. -
14.
LTL:
Explanation:Cannot craft sticks without 2 planks. -
15.
LTL:
Explanation:Cannot smelt iron without a furnace, fuel, and iron ore. -
16.
LTL:
Explanation:Cannot equip a wooden pickaxe unless the agent has one. -
17.
LTL:
Explanation:Cannot equip a stone pickaxe unless the agent has one. -
18.
LTL:
Explanation:Cannot place a crafting table unless the agent has one. -
19.
LTL:
Explanation:Cannot place a furnace unless the agent has one.
The soft LTL rules implemented by the critic are as follows:
1.
LTL:
Explanation:If the agent lacks logs, planks, sticks, and an iron pickaxe, it should mine wood logs. -
2.
LTL:
Explanation:If the agent has logs but no planks, it should craft planks. -
3.
LTL:
Explanation:If the agent has planks but no sticks, it should craft sticks. -
4.
LTL:
Explanation:If the agent has 4 planks and 2 sticks but no crafting table is nearby or already crafted, it must craft a crafting table. -
5.
LTL:
Explanation:If the agent has 4 planks, 2 sticks, and a crafting table but is not near one, it should place the crafting table. -
6.
LTL:
Explanation:If the agent has 3 planks, 2 sticks, is near a crafting table, and doesn’t already have a wooden pickaxe, it should craft one. -
7.
LTL:
Explanation:If the agent has a wooden pickaxe but it isn’t equipped and it lacks three cobblestones, it should equip the pickaxe. -
8.
LTL:
Explanation:If equipped with a wooden pickaxe and lacking stone, the agent should mine cobblestone. -
9.
LTL:
Explanation:If the agent has 3 cobblestones and a wooden pickaxe equipped, it should craft a stone pickaxe. -
10.
LTL:
Explanation:If the agent has 3 iron ore, is near a furnace, has fuel, and lacks 3 ingots, it should smelt iron. -
11.
LTL:
Explanation:If the agent has the ingredients but no iron pickaxe, it should craft one. -
12.
LTL:
Explanation:If the agent owns an iron pickaxe but hasn’t equipped it, it should equip the pickaxe. -
13.
LTL:
Explanation:If an iron pickaxe is equipped and no diamonds are nearby, explore downward. -
14.
LTL:
Explanation:Do not explore for diamonds unless an iron pickaxe is available and equipped. -
15.
LTL:
Explanation:Do not explore if iron and coal are already known to be nearby.
A.3.2 InnerMonologue
Since our safety violations in SayCan simply lead to environmental feedback and new laws, we do not include any hand-engineered safety laws in InnerMonologue.
LTL:
Explanation: If you don’t have logs, planks, or sticks, mine logs. -
LTL:
Explanation: If you have logs but no planks, craft planks. -
LTL:
Explanation: If you have 4 planks, aren’t near a crafting table, and don’t have one, craft a crafting table. -
LTL:
Explanation: If you have a crafting table and a plank, but aren’t near one, place the crafting table. -
LTL:
Explanation: If you have 3 planks, 2 sticks, are near a crafting table, and don’t have a wooden pickaxe, craft one. -
LTL:
Explanation: If you have 3 cobble, 2 sticks, are near a crafting table, and don’t have a stone pickaxe, craft one. -
LTL:
Explanation: If you have 8 cobble, are near a crafting table, and don’t have or see a furnace, craft one. -
LTL:
Explanation: If you have iron ore and fuel, are near a furnace, and don’t already have iron ingots or a pickaxe, smelt iron. -
LTL:
Explanation: If you have 3 iron ingots, 2 sticks, are near a crafting table, and don’t have an iron pickaxe, craft one. -
LTL:
Explanation: If you have an iron pickaxe but it’s not equipped, equip it. -
LTL:
Explanation: If diamonds are nearby and an iron pickaxe is equipped, mine the diamond. -
LTL:
Explanation: If no diamonds are visible and an iron pickaxe is equipped, explore downward for diamonds. -
LTL:
Explanation: If diamonds are visible or no iron pickaxe is equipped, do not explore downward. -
LTL:
Explanation: If no pickaxe is equipped, don’t mine stone. -
LTL:
Explanation: Don’t mine iron ore unless a stone or iron pickaxe is equipped. -
LTL:
Explanation: Don’t craft a furnace without 8 cobblestone. -
LTL:
Explanation: Don’t mine coal without a pickaxe equipped. -
LTL:
Explanation: Don’t craft a wooden pickaxe unless you have the materials and don’t already have one. -
LTL:
Explanation: Don’t craft a stone pickaxe unless you have materials and don’t already have one. -
LTL:
Explanation: Don’t craft an iron pickaxe unless you have materials and don’t already have one.